4. Write an equation of the line parallel to the line y = 2 − 3 x at (4, −2) and sketch its graph.
Answer by Alan3354(69443)

Write an equation of the line parallel to the line y = 2-3x at (4,-2) and sketch its graph.
----------
The slope of the line, m, is -3
Use y = mx + b and the point to find b, the y-intercept
-2 = -3*4 + b
b = 10
-----
--> y = -3x + 10